1. solve each of the following

a) (6xy+2y2-5)dx +(3x2+4xy-6)dy=0 y(1)=2

b)x2d2y/dx2+xdy/dx+4y=2xlnx



1
Expert's answer
2021-07-05T16:54:07-0400

a) "(6xy+2y^2-5)dx +(3x^2+4xy-6) dy=0" , "y(1)=2"


"M(x,y)=6xy+2y^2-5 \implies {\partial M(x,y)\over \partial y}=M_y=6x+4y"


"N(x,y)=3x^2+4xy-6 \implies {\partial N(x,y)\over \partial x}=N_x=6x+4y"


Hence exact since "M_y=N_x"


There exist a function "\mu(x,y)" such that

"{\partial \mu(x,y) \over \partial x}=M(x,y)=6xy+2y^2-5"


"{\partial \mu(x,y) \over \partial y}=N(x,y)=3x^2+4xy-6"


let "\mu(x,y)=\int^xM(x,y)dx" "+g(y)"


"=\int^x(6xy+2y^2-5)dx" "+g(y)"


"=3x^2y+2xy^2-5x" "+g(y)"


"\implies {\partial \mu (x,y) \over \partial y}=3x^2+4xy+{ \partial g(y)\over \partial y}"


"\implies 3x^2+4xy+{ \partial g(y)\over \partial y}=3x^2+4xy-6"


"\implies { \partial g(y)\over \partial y}=-6"


"\implies \int dg(y)=-6 \int dy"


"\implies g(y)=-6y"


"\implies \mu(x,y)=3x^2y+2xy^2-5x-6y"


"\therefore" The general solution is thus "3x^2y+2xy^2-5x-6y=c"

We can find the value of Cc using the initial condition "y(1)=2"

"\implies3(1)^2(2)+2(1)(2)^2-5(1)-6(2)=c \implies c=-3"

"\therefore" the particular solution is

"3x^2y+2xy^2-5x-6y=-3"

Rearranging the equation gives

"3x^2y+2xy^2-5x+3=6y"

dividing through by 6 gives

"y={1\over 2}x^2y+{1\over 3}xy^2-{5\over 6}x+{1\over 2}"

b) "x^2{ d^2y\over dx^2}+x{dy\over dx}+4y=2xlnx"

This is a Cauchy's linear differential equation

let "x=e^z \implies z=lnx"

let "x{dy\over dx}=Dy" , "D={d\over dz}"

let "x^2{ d^2y\over dx^2}=D(D-1)y"

therefore the given equation becomes

"=D(D-1)y+Dy+4y=2ze^z"

"=(D^2-D+D+4)y=2ze^z"

"=(D^2+4)y=2ze^z"

Next we need to find C.F as follows

"D^2+4=0\implies" "D=2i" and "D=-2i" where "i" is an imaginary number

"\implies C.F=C_1e^{-2iz}+C_2e^{2iz}=C_1sin2z+C_2cos2z"

but "z=log" "x"

"\therefore" "C.F=C_1sin 2logx+C_2cos2logx"

Next we need to find "P.I" as follows


"P.I={1\over (D^2+4)}2ze^z=2e^z[{1\over (D+1)^2+4}\times z]=2e^z[{1\over D^2+2D+5}]\times z"

"={2e^z\over 5}[{1\over 1+{D^2+2D\over 5}}]z"

"={2e^z\over 5}[1+{D^2+2D\over 5}]^{-1}z"

"={2e^z\over 5}[1-{D^2+2D\over 5}+...]z"

"={2e^z\over 5}[z-{0^2+2\over 5}]"

"={2e^z\over 5}[z-{2\over 5}]"

"={2ze^z\over 5}-{4e^z\over 25}"

but "x=e^z" and "z=lnx"

"\implies P.I={2xlogx\over 5}-{4logx\over 25}"

"\therefore" The general solution is given as


"y=CF+PI"

"y=C_1sin 2logx+C_2cos2logx+{2xlogx\over 5}-{4logx\over 25}"